Suppose an industry is monopolized, and the demand for the product sold by the firm is given by: Q = 800 ‒ 2P. At what price range should the monopoly firm raise the price in order to increase revenues?

at prices greater than $200
at prices greater than $100
at prices greater than $400
at prices less than $200
at prices less than $100
at prices less than $400
Answer #1

Solution: a price less than $400

Explanation: The demand would be inelastic when price is $400 i.e.

dq/dp = - 2/800P

dq/dp = -1/400 * 400 = -1